OpenStudy (anonymous):

Determine a region whose area is equal to the given limit. Do not evaluate the limit. lim┬(n→∞)⁡∑_(i=1 )^n 2/n (1+2i/n)^18

OpenStudy (anonymous):

can you possibly use the equation thing on the bottom to write this out?

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \sum_{i=1}^{n} \left( 1+ 2i/n\right)^{18}\]

OpenStudy (anonymous):

I know that the interval [a,b] is [1,3] because \[x _{i}=a+i \Lambda x=1+2i/n\]

OpenStudy (anonymous):

I just don't know how to find the region from this information